OpenStudy (anonymous):

Use the discriminant to determine the nature of the roots of x62 – 3x + 3 = 0.

OpenStudy (anonymous):

its ment to say x^2-3x+3=0

OpenStudy (anonymous):

a=1, b=-3, c=3 [D=(-3)^2-4(1)(3)} =9-12=-3<0 So Roots are imaginary.

OpenStudy (anonymous):

THANK YOU SO MUCH